# UniTracker: Learning Universal Whole-Body Motion Tracker for Humanoid Robots

### v2

@@ Abstract
Achieving expressive and generalizable whole-body motion control is essential for deploying humanoid robots in real-world environments. In this work, we propose UniTracker, a three-stage training framework that enables robust and scalable motion tracking across a wide range of human behaviors. In the first stage, we train a teacher policy with privileged observations to generate high-quality actions. In the second stage, we introduce a Conditional Variational Autoencoder (CVAE) to model a universal student policy that can be deployed directly on real hardware. The CVAE structure allows the policy to learn a global latent representation of motion, enhancing generalization to unseen behaviors and addressing the limitations of standard MLP-based policies under partial observations. Unlike pure MLPs that suffer from drift in global attributes like orientation, our CVAE-student policy incorporates global intent during training by aligning a partial-observation prior to the full-observation encoder. In the third stage, we introduce a fast adaptation module that fine-tunes the universal policy on harder motion sequences that are difficult to track directly. This adaptation can be performed both for single sequences and in batch mode, further showcasing the flexibility and scalability of our approach. We evaluate UniTracker in both simulation and real-world settings using a Unitree G1 humanoid, demonstrating strong performance in motion diversity, tracking accuracy, and deployment robustness.
![[Uncaptioned image]](x1.png)
Keywords: Whole-Body Humanoid Robot Tracker, Imitation Learning, Reinforcement Learning
1 Introduction
@@
Recent research has explored various control interfaces tailored to humanoid whole-body controller. These can be broadly categorized into dense and sparse control signals. Dense signals, such as teleoperation [1, 2, 3, 4, 5, 6], offline motion datasets [7, 8, 9, 10, 11, 12, 13], and video-based motion estimation [14, 15], provide rich trajectory-level information. In contrast, sparse signals such as high-level task commands and VR-based guidance [16, 17] offer minimal information and often lead to reduced motion quality. In this work, we focus on universal whole-body motion tracking, where the input is a reference motion sequence and the goal is to track it robustly and expressively using a single policy.
@@
To address the aforementioned limitations of existing teacher-student frameworks, we introduce UniTracker, a unified and expressive whole-body tracking framework that integrates a Conditional Variational Autoencoder (CVAE) [19] into the student policy architecture. By explicitly modeling a structured latent space conditioned on future motion references, UniTracker enables the policy to generate diverse and high-fidelity behaviors even under partial observations. From a probabilistic standpoint, the latent variable captures the inherent ambiguity in the mapping from observations to actions, allowing the policy to model a distribution over plausible motor behaviors rather than collapsing to a single deterministic output. This capability enhances motion expressiveness and significantly improves generalization across diverse and unseen movement patterns.
@@
Our main contributions are summarized as follows: i)A three-stage training framework for universal whole-body tracking: We design a modular pipeline comprising a privileged teacher for data generation, a CVAE-based student policy for deployment under partial observations, and a lightweight fast adaptation phase for motion-specific fine-tuning. ii)Diversity-aware and gloabl context-integrated policy via CVAE modeling: We employ a Conditional Variational Autoencoder to capture motion diversity and encode global context, enabling expressive behaviors and reducing global inconsistencies by aligning a globally-informed encoder with a partial-observation prior. iii)Fast adaptation on top of a universal policy: We introduce a rapid fine-tuning mechanism that adapts the universal policy to challenging or out-of-distribution motions, supporting both single-sequence and batch-level adaptation with minimal overhead. iv)Extensive validation on real-world hardware: We demonstrate that UniTracker robustly tracks over 8,100 human motions on a 29-DoF humanoid using a single policy, outperforming strong teacher-student baselines in accuracy, robustness, and generalization.
2 Method
2.1 Problem Formulation
@@
The remainder of this section is organized as follows. Section 2.2 introduces the construction of a high-quality humanoid motion dataset for policy training. Section 2.3 describes the training of an oracle policy in simulation, which is designed both to maximize the expressiveness of motion tracking. Section 2.4 details the distillation of the oracle policy into a deployable student using a CVAE-based framework. Section 2.5 presents our fast adaptation strategy, which enables rapid fine-tuning based on the universal tracker. An overview of the proposed universal whole-body motion tracking framework is provided in Figure 2.
2.2 Humanoid Motion Dataset Curation
A large-scale humanoid motion dataset serves as fuel for training a universal motion tracker. Our dataset is primarily derived from the publicly available AMASS [22] dataset, filtered to exclude interactions and sequences shorter than 10 frames. This yields a final training set of 11,313 human motions, represented using the SMPL [23] parameters. The SMPL model parameterizes the human body through shape parameters , pose parameters and root translation . denotes the SMPL function, where maps the SMPL parameters to the positions of vertices of a triangular human mesh.
@@ 2.2 Humanoid Motion Dataset Curation
2.3 Oracle Policy Training in Simulation
@@
Reward Design. We formulate the reward as a weighted sum of three components: 1) task rewards for motion tracking, 2)regularization, and 3) penalty, detailed in Table 3. We apply curriculum learning to the regularization terms and penalty terms such that the policy could better focus on the motion tracking task itself and gradually take the penalty and regularization into account for more reasonable behaviors.
@@
Domain Randomization. Domain randomization has been proved to be a critical source of robustness and generalization for successful sim-to-real transfer [11, 4, 6]. Existing domain randomization terms can be broadly categorized into two classes: asset property randomization and environmental dynamic randomization. Asset property randomization, focusing on parameters such as friction coefficients, center of mass offsets, and link masses, aims to prevent the policy from overfitting to a specific asset configuration. In contrast, environmental dynamic randomization, focusing on PD gains, applied torques, or other external perturbations (e.g., pushing the robot), introduces dynamic variations to simulate the environmental uncertainties encountered during real-world deployment. While previous works typically apply both forms of randomization concurrently throughout training, we propose decoupling them to better balance the tradeoff between motion tracking expressiveness and robustness. Hence, in the first stage, we exclusively randomize asset properties to maximize the policy’s ability to accurately track diverse motions. This approach not only builds an upper bound for subsequent policies but also yields a proxy agent capable of generating high-quality action signals in simulators.
2.4 Hierarchical Controller via Online Distillation
Deployable State Space Design. Since certain privileged information in the oracle state space is unavailable in real-world deployment, we define a deployable state space based on data accessible on humanoid robots. The proprioception is defined as where and denote joint positions and velocities, refers to the root angular velocity, is the gravity vector and is the previous action. These terms are stacked over the past 25 steps to form the proprioceptive input. The goal state is defined as , where is the reference pose height, and are the reference and current root orientations, and are the reference root linear and angular velocities, is the current root angular velocity, and are the reference rigid body and root positions. The first four terms are rotated into the current pose’s local coordinate while the last term is rotated into the reference pose’s local frame.
where and is the KL divergence between the prior and encoder.
2.5 Fast Adaption on top of a Universal Policy
To address these cases, we introduce a lightweight residual decoder appended to the universal motion tracking framework, to enable fast adaptation to challenging motions. During training, we initialize the model with the universal policy checkpoint and freeze all existing parameters. Only the residual decoder is fine-tuned using reinforcement learning, the reward is the same as Section 2.3. The input to the residual decoder is identical to that of the prior network. We directly use explicit motion references instead of latent representations for two main reasons. First, such out-of-distribution motions may not be adequately captured by the learned CVAE prior. Second, the goal at this stage is to quickly overfit to specific motion sequences rather than to maintain generalization. The final action applied to the environment is obtained by summing the output of the universal policy and the residual correction:
SR↑ | MPKPE↓ | Vel-Dist↓ | Acc-Dist↓ | MPKPE↓ | Vel-Dist↓ | Acc-Dist↓ | |
(a) Compare with Baselines | |||||||
Dagger without CVAE | 88.21 | 84.79 | 5.60 | 2.97 | 84.70 | 5.03 | 2.41 |
Train from Scratch | 72.11 | 103.94 | 8.78 | 6.33 | 98.61 | 8.47 | 6.31 |
Ours | 91.83 | 82.62 | 4.27 | 1.83 | 78.95 | 3.65 | 1.39 |
(b) Ablation with Architecture Design | |||||||
Actor with Explicit Reference | 88.29 | 85.31 | 5.76 | 3.03 | 84.96 | 4.58 | 2.15 |
Actor without Explicit Reference | 91.83 | 82.62 | 4.27 | 1.83 | 78.95 | 3.65 | 1.39 |
(c) Ablation with KL Residual | |||||||
KL without Residual | 85.49 | 91.24 | 6.02 | 4.68 | 90.79 | 6.01 | 4.23 |
KL with Residual | 91.83 | 82.62 | 4.27 | 1.83 | 78.95 | 3.65 | 1.39 |
(d) Ablation with KL Coefficient | |||||||
KL Coef = 1.0 | 81.18 | 97.23 | 7.75 | 6.04 | 96.53 | 5.39 | 2.41 |
KL Coef = 0.1 | 91.83 | 82.62 | 4.27 | 1.83 | 78.95 | 3.65 | 1.39 |
KL Coef = 0.01 | 86.10 | 89.88 | 7.99 | 5.47 | 88.01 | 5.23 | 2.37 |
KL Coef = 0.001 | 80.55 | 105.56 | 9.21 | 7.98 | 104.63 | 5.38 | 2.40 |
(e) Ablation with Future Window Size | |||||||
Window Size = 1 | 90.77 | 85.98 | 5.20 | 3.12 | 85.00 | 4.79 | 2.91 |
Window Size = 5 | 91.83 | 82.62 | 4.27 | 1.83 | 78.95 | 3.65 | 1.39 |
Window Size = 10 | 91.28 | 82.14 | 4.77 | 1.95 | 78.78 | 3.90 | 1.87 |
Window Size = 20 | 90.29 | 82.96 | 5.01 | 1.78 | 80.24 | 4.33 | 1.48 |
(f) Ablation with Latent Dimension | |||||||
Latent Dimension = 32 | 88.82 | 87.25 | 4.79 | 2.36 | 86.52 | 4.67 | 2.21 |
Latent Dimension = 64 | 91.83 | 82.62 | 4.27 | 1.83 | 78.95 | 3.65 | 1.39 |
Latent Dimension = 128 | 90.54 | 83.75 | 4.46 | 2.14 | 82.92 | 4.45 | 2.08 |
Latent Dimension = 256 | 90.72 | 83.99 | 4.21 | 1.97 | 83.71 | 4.19 | 1.77 |
(g) Analysis of CVAE diversity | |||||||
Decoder with Stochastic Latent | 91.77 | 83.21 | 4.59 | 1.82 | 79.07 | 3.72 | 1.33 |
Decoder with Deterministic Latent | 91.83 | 82.62 | 4.27 | 1.83 | 78.95 | 3.65 | 1.39 |
3 Experimental Results
3.1 Experiment Setup
We evaluate UniTracker in both simulated and real-world environments. In simulation, policies are trained using IsaacGym [26] with 8192 parallel environments under domain randomization [27]. The training data is derived from the AMASS dataset [22] and filtered by PHC [25]. We further validate our approach through sim-to-sim transfer by deploying the trained policies in MuJoCo, and conduct extensive ablation and comparative studies. Performance is assessed using four key metrics: Success Rate (SR), Mean Per Keypoint Position Error (MPKPE), Velocity Distance(Vel-Dist) and Acceleration Distance(Acc-Dist). SR reflects the overall viability and stability of the policy; MPKPE respectively quantify the accuracy of keypoint tracking in the world coordinate frame; Vel-Dist quantifies the difference in joint velocities between the reference motion and the executed motion of the robot; Acc-Dist quantifies the difference in joint acceleration between the reference motion and the executed motion of the robot. For real-world evaluation, we deploy UniTracker on the Unitree G1 humanoid robot, which stands 1.3 meters tall and has 29 degrees of freedom. We control 23 DoFs by locking the 6 wrist joints.
3.2 Baselines
What kind of motion tracker yields the best tracking performance? We include two baselines: (1) a universal policy trained from scratch without the teacher-student architecture, and (2) a teacher-student framework using DAgger in the second stage without a CVAE. Direct comparisons with prior methods are not included, as their implementations are not publicly available. Experimental results in Table 1 show that UniTracker consistently outperforms our baselines across all evaluation metrics.
3.3 Ablation Studies
What model architecture best balances diversity, expressiveness and robustness? To answer this question, we first investigated whether the RL actor (the decoder in the CVAE) should explicitly take the reference motion as part of its input. We evaluated two variants: one where the actor input is and another where the reference motion is excluded. The results show that when the actor receives the reference motion directly, the influence of the latent variable vanishes. In this case, the behavior of the model closely resembles that of a standard DAgger setup, as the strong reference input causes the policy to ignore the latent guidance. We further evaluated a pure DAgger setup with an MLP-based actor in the second stage. While this design performs well on motions seen during training, it degrades noticeably on out-of-distribution (OOD) motions. The strong reliance on explicit reference inputs reduces the policy’s robustness and its ability to generalize to unseen or novel motion sequences.
In Figure 3, we evaluate the generalization capability of UniTracker in MuJoCo by testing it on a challenging lateral squat motion, which demands strong balance control from the robot. Although UniTracker cannot precisely track all keypoints of the reference motion, it successfully produces a plausible lateral squat behavior while maintaining balance. In contrast, the DAgger baseline without the CVAE fails to remain stable and results in the robot falling, highlighting the importance of latent motion modeling for generalization to unseen motions. We have uploaded more real-world OOD videos on our website, showcasing the generalization ability of our policy.
In Table 2, we evaluate the robustness of our policy by incrementally adding noise to the observations. In this table, a noise level of 0 indicates no observation noise, while higher levels correspond to increasing noise intensity. We observe that UniTracker consistently outperforms the variant without CVAE in both Success Rate (SR) and Mean Per Keypoint Position Error (MPkPE). Moreover, UniTracker exhibits a slower degradation in performance as noise increases, indicating stronger robustness under observation perturbations.
What hyperparameter choices lead to the best performance in whole-body motion tracking? We conduct ablation studies on three key hyperparameters in our framework: the CVAE latent dimension, the future window size of the reference motion used in the CVAE prior, and the weight of the KL loss term. For the latent dimension, we experiment with values of 32, 64, 128, and 256. We observe that a dimension of 64 yields the best performance. Larger latent sizes make the CVAE harder to train and do not lead to further improvements. Results are shown in Table 1 For the future window, which controls how many future frames of reference motion are included in the prior input, we test window sizes of 1, 5, 10, and 20 frames. The results indicate that using 5 future frames provides the best trade-off between responsiveness and stability. Lastly, we tune the KL loss coefficient with values of 1.0, 0.1, 0.01, and 0.001. We find that setting the weight to 0.1 leads to the best policy performance, effectively balancing latent regularization and reconstruction quality.
3.4 Downstream Applications
We evaluate two downstream applications of our framework: motion generation and video-based motion estimation. For motion generation, we use the MDM [28] model to produce SMPL-based [23] motion sequences conditioned on text input. These motions are then retargeted to the G1 humanoid robot for tracking. Both MuJoCo [29] simulation and real-world results demonstrate that our policy can accurately track motions generated from text prompts. For video estimation, we record a long motion sequence using a monocular camera, and convert it to SMPL format using GVHMR [30]. The resulting motion is retargeted to the G1 robot. Experiments in both MuJoCo and the real world show that our policy successfully tracks the reconstructed motions. These two applications highlight the strong generalization capability of our policy across different types of reference inputs.
3.5 Light Weight Fast Adaption
We compared our fast adaptation module against a training-from-scratch baseline. In the latter, the observations of our model remain consistent with those of the second-stage student policy, and the reward function is identical to that used in the fast adaptation module. For evaluation, we selected three challenging motion sequences—Round-Horse Kick, Side Kick, and Martial Art—along with the entire AMASS test dataset. As shown in Figure 6, our fast adaptation approach converges substantially faster than training from scratch in terms of both episode length and cumulative reward. These results underscore the advantages of our second-stage universal policy and highlight the effectiveness of the pretrained model in enabling rapid adaptation.
Methods | All AMASS Train Dataset | |
SR | MPKPE | |
(a) Noise Level 0 | ||
Dagger without CVAE | 88.21 | 84.79 |
Ours | 91.83 | 82.62 |
(b) Noise Level 1 | ||
Dagger without CVAE | 85.79 | 88.65 |
Ours | 90.26 | 83.84 |
(c) Noise Level 2 | ||
Dagger without CVAE | 79.58 | 93.71 |
Ours | 86.79 | 87.31 |
4 Related Work
4.1 Whole-Body Controller for Humanoid Robots
Whole-body control is essential for enabling humanoid robots to perform a wide variety of complex tasks. Prior to the rise of reinforcement learning, researchers primarily relied on traditional optimization-based control methods for humanoid whole-body control [31, 32, 33, 34, 35, 36, 37]. These approaches typically required explicit mathematical modeling of both the robot and its environment, followed by real-time optimization to compute the robot’s next action. However, such methods often struggle to adapt to environmental variations, resulting in limited robustness. Additionally, they impose heavy computational demands during online execution.
To overcome these limitations, reinforcement learning (RL) has emerged as a powerful alternative, offering the ability to learn adaptive, robust control policies directly from interaction with the environment without relying on explicit modeling. Current reinforcement learning–based whole-body controllers for humanoid robots can be categorized by the source of their control signals, including teleoperation [1, 2, 3, 4, 5, 6], offline motion datasets [7, 8, 10, 11], video-based motion estimation [14, 15] [5, 6], and high-level task commands [16, 17]. Teleoperation involves directly controlling the humanoid robot in real time using human input, often through motion capture systems or wearable sensors, allowing the robot to mimic human movements with high fidelity. Representative works in this area include Twist [1] and H2O [4], both of which adopt a two-stage teacher-student framework. The primary difference lies in the design of the policy’s observation space. Offline motion datasets consist of pre-collected human or humanoid motion sequences, which are used as references for training control policies through motion imitation. Representative works are Exbody [11] and Exbody2 [10], which begins by carefully curating an offline motion dataset and then decouples upperand lower-body motions as much as possible, aiming to maintain stability in the lower body while encourage diversity and expressiveness in the upper body. In addition, the most recent work, GMT [12], is the first to demonstrate tracking of 8,000 motions using a single unified policy. Video-based motion estimation methods leverage visual input from videos to extract human motion data, which can then be used to guide humanoid robot control policies. This approach enables learning from large-scale, diverse motion sources without requiring direct human demonstration. A representative work is VideoMimic [14], which develops a real-to-sim-to-real pipeline to model both the robot and its surrounding environment. Task commands refer to high-level, sparse control signals that specify desired outcomes or goals, such as walking direction or target position, rather than detailed joint-level motions, enabling efficient whole-body control through abstraction. Representative works include Hover [16]and HugWBC [17]. Hover unifies multiple control modes into a single policy, enabling seamless transitions while retaining the strengths of each mode, thus providing a robust and scalable humanoid control solution. HugWBC designs a general task and behavior command space and employs techniques such as symmetrical loss and intervention training. This enables real-world humanoid robots to perform a variety of natural gaits—including walking, jumping, and hopping.
@@ 4.1 Whole-Body Controller for Humanoid Robots
5 Conclusion

### v3

@@ Abstract
Achieving generalizable whole-body motion control is essential for deploying humanoid robots in real-world environments. However, existing MLP-based policies trained under partial observations often suffer from limited expressiveness and struggle to maintain global consistency. These shortcomings manifest as less expressive motion, orientation drift, and poor generalization across diverse behaviors. To address these limitations, we propose UniTracker, a three-stage framework for scalable and adaptive motion tracking. The first stage learns a privileged teacher policy that produces high-fidelity reference actions. Building on this, the second stage trains a CVAE-based universal policy that captures a global latent representation of motion, enabling robust performance under partial observations. Crucially, we align the partial-observation prior with a full-observation encoder, injecting global intent into the latent space. In the final stage, a lightweight adaptation module fine-tunes the student policy on challenging sequences, supporting both per-instance and batch adaptation. We validate UniTracker in simulation and on a Unitree G1 humanoid robot, demonstrating superior tracking accuracy, motion diversity, and deployment robustness compared to current baselines.
I INTRODUCTION
@@
Recent research has explored various control interfaces tailored to humanoid whole-body controller. These can be broadly categorized into dense and sparse control signals. Dense signals, such as teleoperation [1, 2, 3, 4, 5, 6], offline motion datasets [7, 8, 9, 10, 11, 12, 13], and video-based motion estimation [14, 15], provide rich trajectory-level information. In contrast, sparse signals such as high-level task commands and VR-based guidance [16, 17] offer minimal information and often lead to reduced motion quality. In this work, we focus on universal whole-body motion tracking, where the input is reference motion sequences and the goal is to track them robustly and expressively using a single policy.
@@
To address the aforementioned limitations of existing teacher-student frameworks, we introduce UniTracker, a unified and expressive whole-body tracking framework that integrates a Conditional Variational Autoencoder (CVAE) [19] into the student policy architecture. By explicitly modeling a structured latent space conditioned on future motion references, UniTracker enables the policy to generate diverse and high-fidelity behaviors even under partial observations. From a probabilistic standpoint, the latent variable captures the inherent ambiguity in the mapping from observations to actions, allowing the policy to model a distribution over plausible motor behaviors rather than collapsing to a single deterministic output. This capability enhances motion expressiveness and significantly improves generalization across diverse and unseen motion patterns.
@@
Our main contributions are summarized as follows: i) A three-stage training framework for universal whole-body tracking: We design a modular pipeline comprising a privileged teacher for data generation, a CVAE-based student policy for deployment under partial observations, and a lightweight fast adaptation phase for motion-specific fine-tuning. ii) Diversity-aware and global context-integrated policy via CVAE modeling: We employ a Conditional Variational Autoencoder to capture motion diversity and encode global context, enabling expressive behaviors and reducing global inconsistencies by aligning a globally-informed encoder with a partial-observation prior. iii) Fast adaptation on top of a universal policy: We introduce a rapid fine-tuning mechanism that adapts the universal policy to challenging or out-of-distribution motions, supporting both single-sequence and batch-level adaptation with minimal overhead. iv) Extensive validation on real-world hardware: We demonstrate that UniTracker robustly tracks over 8,100 human motions on a 29-DoF humanoid using a single policy, outperforming strong teacher-student baselines in accuracy, robustness, and generalization.
II METHOD
II-A Problem Formulation
@@
The remainder of this section is organized as follows. Section II-B introduces the construction of a high-quality humanoid motion dataset for policy training. Section II-C describes the training of an oracle policy in simulation, which is designed both to maximize the expressiveness of motion tracking. Section II-D details the distillation of the oracle policy into a deployable student using a CVAE-based framework. Section II-E presents our fast adaptation strategy, which enables rapid fine-tuning based on the universal tracker. An overview of the proposed universal whole-body motion tracking framework is provided in Figure 2.
II-B Humanoid Motion Dataset Curation
A large-scale humanoid motion dataset serves as fuel for training a universal motion tracker. Our dataset is primarily derived from the publicly available AMASS dataset [22]. We first exclude sequences involving human-object interactions and discard clips shorter than 10 frames. We then apply the filtering protocol introduced in PHC [23] to further refine the dataset, ensuring high-quality motion sequences suitable for whole-body tracking. This yields a final training set of 8179 human motions, represented using the SMPL [24] parameters. The SMPL model parameterizes the human body through shape parameters , pose parameters and root translation . denotes the SMPL function, where maps the SMPL parameters to the positions of vertices of a triangular human mesh.
@@ II-B Humanoid Motion Dataset Curation
II-C Oracle Policy Training in Simulation
@@
Reward Design. We formulate the reward as a weighted sum of three components: 1) task rewards for motion tracking, 2)regularization, and 3) penalty, detailed in Table I. We apply curriculum learning to the regularization terms and penalty terms such that the policy could better focus on the motion tracking task itself and gradually take the penalty and regularization into account for more reasonable behaviors.
@@
Domain Randomization. Domain randomization is a key technique for improving sim-to-real robustness [11, 4, 6]. Existing strategies generally fall into two categories: asset property randomization and environmental dynamics randomization. The former modifies physical parameters such as friction, center of mass, and link mass to prevent overfitting to a specific asset configuration. The latter perturbs dynamic elements such as PD gains, torque noise, or external forces to simulate real-world uncertainty. While prior work often applies both types concurrently, we decouple them to better balance tracking expressiveness and robustness. In the first stage, we randomize only asset properties, allowing the teacher policy to focus on high-fidelity motion tracking. This design provides a strong performance upper bound and serves as a proxy agent for generating high-quality reference actions.
II-D Hierarchical Controller via Online Distillation
Deployable State Space Design.
@@ II-D Hierarchical Controller via Online Distillation
Since certain privileged information in the oracle state space is unavailable in real-world deployment, we define a deployable state space based on data accessible on humanoid robots. The proprioception is defined as where and denote joint positions and velocities, refers to the root angular velocity, is the gravity vector and is the previous action. These terms are stacked over the past 25 steps to form the proprioceptive input. The goal state is defined as , where is the reference pose height, and are the reference and current root orientations, and are the reference root linear and angular velocities, is the current root angular velocity, and are the reference rigid body and root positions. These terms are stacked over the future 5 steps. The first four terms are rotated into the current pose’s local coordinate while the last term is rotated into the reference pose’s local frame.
SR↑ | MPKPE↓ | Vel-Dist↓ | Acc-Dist↓ | MPKPE↓ | Vel-Dist↓ | Acc-Dist↓ | |
(a) Compare with Baselines | |||||||
OmniH2O [6] | |||||||
Exbody2 [10] | |||||||
Dagger without CVAE | |||||||
Train from Scratch | |||||||
Ours | |||||||
(b) Ablation with Architecture Design | |||||||
Actor with Explicit Reference | |||||||
Actor without Explicit Reference | |||||||
(c) Ablation with KL Residual | |||||||
KL without Residual | |||||||
KL with Residual | |||||||
(d) Ablation with KL Coefficient | |||||||
KL Coef = 1.0 | |||||||
KL Coef = 0.1 | |||||||
KL Coef = 0.01 | |||||||
KL Coef = 0.001 | |||||||
(e) Ablation with Future Window Size | |||||||
Window Size = 1 | |||||||
Window Size = 5 | |||||||
Window Size = 10 | |||||||
Window Size = 20 | |||||||
(f) Ablation with Latent Dimension | |||||||
Latent Dimension = 32 | |||||||
Latent Dimension = 64 | |||||||
Latent Dimension = 128 | |||||||
Latent Dimension = 256 |
where and is the KL divergence between the prior and encoder.
II-E Fast Adaption on top of a Universal Policy
To address these cases, we introduce a lightweight residual decoder appended to the universal motion tracking framework, to enable fast adaptation to challenging motions. During training, we initialize the model with the universal policy checkpoint and freeze all existing parameters. Only the residual decoder is fine-tuned using reinforcement learning, the reward is the same as Section II-C. The input to the residual decoder is identical to that of the prior network. We directly use explicit motion references instead of latent representations for two main reasons. First, such out-of-distribution motions may not be adequately captured by the learned CVAE prior. Second, the goal at this stage is to quickly overfit to specific motion sequences rather than to maintain generalization. The final action applied to the environment is obtained by summing the output of the universal policy and the residual correction:
III Experimental Results
III-A Experiment Setup
We evaluate UniTracker in both simulated and real-world environments. In simulation, policies are trained using IsaacGym [26] with 8192 parallel environments under domain randomization [27]. The training data is derived from the AMASS dataset [22] and filtered by PHC [23]. We further validate our approach through sim-to-sim transfer by deploying the trained policies in MuJoCo, and conduct extensive ablation and comparative studies. Performance is assessed using four key metrics: Success Rate (SR), Mean Per Keypoint Position Error (MPKPE), Velocity Distance(Vel-Dist) and Acceleration Distance(Acc-Dist). SR reflects the overall viability and stability of the policy; MPKPE quantify the accuracy of keypoint tracking in the world coordinate frame; Vel-Dist quantifies the difference in joint velocities between the reference motion and the executed motion of the robot; Acc-Dist quantifies the difference in joint acceleration between the reference motion and the executed motion of the robot. For real-world evaluation, we deploy UniTracker on the Unitree G1 humanoid robot, which stands 1.3 meters tall and has 29 degrees of freedom. We control 23 DoFs by locking the 6 wrist joints.
III-B Baselines
What kind of motion tracker yields the best tracking performance? We include two internal baselines to assess the effectiveness of each component in our framework: (1) a universal policy trained from scratch without leveraging the teacher-student architecture; (2) a teacher-student framework using DAgger [18] in the second stage, but without incorporating CVAE-based modeling. In addition, we compare UniTracker against recent state-of-the-art approaches for humanoid motion tracking, including OmniH2O[6] and Exbody2[10]. Experimental results presented in Table II demonstrate that UniTracker consistently outperforms all baselines and prior methods across all evaluation metrics.
III-C Ablation Studies
What model architecture best balances diversity, expressiveness and robustness? To answer this question, we first investigated whether the RL actor (the decoder in the CVAE) should explicitly take the reference motion as part of its input. We evaluated two variants: one where the actor input is and another where the reference motion is excluded. The results show that when the actor receives the reference motion directly, the influence of the latent variable vanishes. In this case, the behavior of the model closely resembles that of a standard DAgger setup, as the strong reference input causes the policy to ignore the latent guidance.
In Figure 4, we evaluate the generalization capability of UniTracker by test the policy in MuJoCo on a challenging crescent kick motion that is unseen during training. We observe that UniTracker successfully tracks this previously unseen behavior, while the variant without CVAE modeling fails to reproduce the motion effectively. This result highlights the importance of structured latent motion modeling for generalization to out-of-distribution motions. Additional real-world out-of-distribution (OOD) examples are provided in the supplementary video, further demonstrating the robustness and adaptability of our approach under novel motion inputs.
In Table III, we evaluate the robustness of our policy by incrementally adding noise to the observations. In this table, a noise level of 0 indicates no observation noise, while higher levels correspond to increasing noise intensity. We observe that UniTracker consistently outperforms the variant without CVAE in both Success Rate (SR) and Mean Per Keypoint Position Error (MPkPE). Moreover, UniTracker exhibits a slower degradation in performance as noise increases, indicating stronger robustness under observation perturbations.
What hyperparameter choices lead to the best performance in whole-body motion tracking? We conduct ablation studies on three key hyperparameters in our framework: the CVAE latent dimension, the future window size of the reference motion used in the CVAE prior, and the weight of the KL loss term. For the latent dimension, we experiment with values of 32, 64, 128, and 256. We observe that a dimension of 64 yields the best performance. Larger latent sizes make the CVAE harder to train and do not lead to further improvements. Results are shown in Table II For the future window, which controls how many future frames of reference motion are included in the prior input, we test window sizes of 1, 5, 10, and 20 frames. The results indicate that using 5 future frames provides the best trade-off between responsiveness and stability. Lastly, we tune the KL loss coefficient with values of 1.0, 0.1, 0.01, and 0.001. We find that setting the weight to 0.1 leads to the best policy performance, effectively balancing latent regularization and reconstruction quality.
III-D Downstream Applications
We evaluate two downstream applications of our framework: motion generation and video-based motion estimation. For motion generation, we use the MDM [28] model to produce SMPL-based [24] motion sequences conditioned on text input. These motions are then retargeted to the G1 humanoid robot for tracking. Both MuJoCo [29] simulation and real-world results demonstrate that our policy can accurately track motions generated from text prompts. For video estimation, we record a motion sequence using a monocular camera, and convert it to SMPL format using GVHMR [30]. The resulting motion is retargeted to the G1 robot. Experiments in both MuJoCo and the real world show that our policy successfully tracks the reconstructed motions. These two applications highlight the strong generalization capability of our policy across different types of reference inputs.
III-E Light Weight Fast Adaption
We compared our fast adaptation module against a training-from-scratch baseline. In the latter, the observations of our model remain consistent with those of the second-stage student policy, and the reward function is identical to that used in the fast adaptation module. For evaluation, we selected one challenging motion sequence along with the entire AMASS test dataset. As shown in Figure 5, our fast adaptation approach converges substantially faster than training from scratch in terms of both episode length and cumulative reward. These results underscore the advantages of our second-stage universal policy and highlight the effectiveness of the pretrained model in enabling rapid adaptation.
Methods | All AMASS Train Dataset | |
SR | MPKPE | |
(a) Noise Level 0 | ||
Dagger without CVAE | ||
Ours | ||
(b) Noise Level 1 | ||
Dagger without CVAE | ||
Ours | ||
(c) Noise Level 2 | ||
Dagger without CVAE | ||
Ours |
IV Related Work
IV-A Whole-Body Controller for Humanoid Robots
Whole-body control is essential for enabling humanoid robots to perform a wide variety of complex tasks. Prior to the rise of reinforcement learning, researchers primarily relied on traditional optimization-based control methods for humanoid whole-body control [31, 32, 33]. These approaches typically required explicit mathematical modeling of both the robot and its environment, followed by real-time optimization to compute the robot’s next action. However, such methods often struggle to adapt to environmental variations, resulting in limited robustness. Additionally, they impose heavy computational demands during online execution.
To overcome these limitations, reinforcement learning (RL) has emerged as a powerful alternative, offering the ability to learn adaptive, robust control policies directly from interaction with the environment without relying on explicit modeling. Current reinforcement learning–based whole-body controllers for humanoid robots can be categorized by the source of their control signals, including teleoperation [1, 2, 3, 4, 5, 6], offline motion datasets [7, 8, 13, 34, 10, 11], video-based motion estimation [14, 15] [5, 6], and high-level task commands [16, 17]. Teleoperation involves directly controlling the humanoid robot in real time using human input, often through motion capture systems or wearable sensors, allowing the robot to mimic human movements with high fidelity. Representative works in this area include Twist [1] and H2O [4], both of which adopt a two-stage teacher-student framework. The primary difference lies in the design of the policy’s observation space. Offline motion datasets consist of pre-collected human or humanoid motion sequences, which are used as references for training control policies through motion imitation. Representative works are Exbody [11] and Exbody2 [10], which begins by carefully curating an offline motion dataset and then decouples upperand lower-body motions as much as possible, aiming to maintain stability in the lower body while encourage diversity and expressiveness in the upper body. In addition, the most recent work, GMT [12], is the first to demonstrate tracking of 8,000 motions using a single unified policy. Video-based motion estimation methods leverage visual input from videos to extract human motion data, which can then be used to guide humanoid robot control policies. This approach enables learning from large-scale, diverse motion sources without requiring direct human demonstration. A representative work is VideoMimic [14], which develops a real-to-sim-to-real pipeline to model both the robot and its surrounding environment. Task commands refer to high-level, sparse control signals that specify desired outcomes or goals, such as walking direction or target position, rather than detailed joint-level motions, enabling efficient whole-body control through abstraction. Representative works include Hover [16]and HugWBC [17]. Hover unifies multiple control modes into a single policy, enabling seamless transitions while retaining the strengths of each mode, thus providing a robust and scalable humanoid control solution. HugWBC designs a general task and behavior command space and employs techniques such as symmetrical loss and intervention training. This enables real-world humanoid robots to perform a variety of natural gaits—including walking, jumping, and hopping.
V Conclusion